Grade 12th passPhysical Chemistry

Potential energy is -27.2 eV in second orbit of He+ then calculate the double of total energy in first exicted state of hydrogen atom?
(Ans is given as -6.8eV)

Radhika
8 Years agoGrade 12th pass

1 Answer

Satwik Patial
7 Years ago
p.E=-2En so En of lithium= -27.2÷2=-13.6eV
So for second orbit n=4 of helium atom z=2 
Or for first excit state n= 2  of hydrogen atom n=1  From formula Enof lithium÷ En of hydrogen = (E0÷Eo)×(zsquare÷nsqare)of litium×(nsquare÷zsquare)of hydrogen =-3.4eV
Double total energy=2En of hydrogen=-6.9eV